OpenStudy (anonymous):

Find the midpoint of the line segment with endpoints (-6,9) and (4,17).

OpenStudy (anonymous):

use the formula \[d = \sqrt{(x _{2}-x _{1})+(y _{2}-y _{1})}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

scratch that, my mistake

OpenStudy (anonymous):

take the average of the x values and y values

OpenStudy (anonymous):

(-6+4)/2 = -1 (9+17)/2 = 13 so the midpoint should be (-1,13)

OpenStudy (anonymous):

I think you use the midpoint formula which is x1+x2/2, y1+y2/2

OpenStudy (amistre64):

just add them together and divide in half

OpenStudy (amistre64):

(-6,9) (4,17) ------ (-2,26)/2 = -1,13